OpenStudy (anonymous):

At the city museum, child admission is $5.40 and adult admission is $9.50 . On Tuesday, tickets 145 were sold for a total sales of $992.10 . How many adult tickets were sold that day?

OpenStudy (anonymous):

@eliassaab @RadEn ?

OpenStudy (anonymous):

@Zale101 ?

OpenStudy (raden):

well, according information above. we can form 2 equations : x + y = 45 5.4x + 9.5y = 992.10 solve for y with y for adult ticket, and x for child ticket

OpenStudy (raden):

by using elimination method is works

OpenStudy (raden):

sorry, the first equation should be x+y=145 not x+y=45 now times 5.4 the 1st equation, get 5.4x + 5.4y = 783 now subtracts both equations : 5.4x + 9.5y = 992.10 5.4x + 5.4y = 783 ------------------- (-) 4.1 y = 209.1 y = 749.1/4.1 = 51 so, there are 51 for adult tickets were sold